# Elastic Constants and Bending Rigidities from Long-Wavelength Perturbation Expansions

### v1

@@ IV.1 Dipole-free materials
Our first example is silicon which is a simple covalent and infrared-inactive solid. The optimized lattice constant of its diamond-like structure from DFT is 5.469 Å, in close agreement with the experimental value 5.431 Å [130]. The Fig. 1 (a) shows the phonon dispersion of silicon with increasing supercell size of IFCs from $4\times 4\times 4$ to $14\times 14\times 14$. Although there is no LO-TO splitting due to its infrared-inactive nature and vanishing Born effective charges, the phonon dispersion only converges from a supercell size of $8\times 8\times 8$. As demonstrated in the inset of Fig. 1 (a), this situation is more severe for long-wavelength acoustic branches around the Brillouin zone center, which is closely linked to the macroscopic elastic properties. The lowest order in the multipole expansion is the quadrupole term, and the QQ term is the only long-range contribution to the dynamical matrix up to the second order in $\mathbf{q}$ in the expansion of Eq. (42). The dynamical quadrupole tensor of silicon can be expressed as $Q_{\kappa\alpha}^{\beta\gamma}=(-1)^{\kappa+1}Q|\varepsilon_{\alpha\beta\gamma}|$ with $\varepsilon_{\alpha\beta\gamma}$ being the Levi-Civita symbol, so it only has one independent component when $\alpha\neq\beta\neq\gamma$. Our fitting procedure based on the Taylor expansion of unscreened charge density response in Eq. (43) gives a value of 13.76 $e\cdot\rm{bohr}$ for $|Q_{x}^{yz}|$ which agrees well with the previous calculations either from DFPT (13.67 $e\cdot\rm{bohr}$ [125]) or a fit to the analytic long-range electron-phonon matrix elements (13.66 $e\cdot\rm{bohr}$ [131]). However, our further interpolation of dynamical matrix with the QQ term taken into account shows an almost unchanged phonon dispersion for silicon, except for the case using a $4\times 4\times 4$ supercell which is too small to ensure the good decay of IFCs in real space. Therefore, the inclusion of quadrupole term does not help to improve the convergence rate of lattice-dynamical properties for silicon, and only the phonon dispersion based on the standard interpolation scheme is illustrated in Fig. 1 (a). We report the dynamical multipole tensors and dielectric properties of silicon in Table 1, although the octupole tensor is not used here due to the vanishing DO interactions in silicon.
@@
Silicon has the highest cubic symmetry and there is only three independent components ($C_{11}$, $C_{12}$ and $C_{44}$) for its elastic tensor (see Sec. II of the SM [132]). The calculated elastic stiffness tensor based on the real space IFCs and long-wavelength expansion as a function of supercell size is shown in Fig. 1 (b). Similar to the results of phonon dispersion, the elastic constants of silicon also begin to converge after a supercell of $8\times 8\times 8$ with the relative error of $4.3\%$ and $2.9\%$ to the reference values from finite-difference calculations for $C_{12}$ and $C_{44}$, respectively, while $C_{11}$ component has already converged after a $6\times 6\times 6$ phonon grid. The results displayed here are calculated based on the total IFCs as silicon is infrared-inactive, without the removal of any high-order multipolar interaction. A further subtraction of the QQ term in the dynamical matrix of silicon does not modify its elastic tensor, which implies that the quadrupole will not couple with the mechanical field therein; this observation is in agreement with the fact that silicon is not a piezoelectric crystal because of the cubic symmetry and the existence of spatial inversion. Using the obtained quadrupole values, the calculated piezoelectric tensor via Martin’s formula [133] is indeed zero. In addition, the detailed comparison of the elastic constants, moduli and sound velocities of silicon calculated from our method based on a supercell of $14\times 14\times 14$ and the finite difference implemented in the thermo_pw code [55] as well as the experimental values [126, 127] are listed in Table 2. The Voigt-Reuss-Hill (VRH) scheme of polycrystalline average is adopted to calculate the bulk ($K$) and shear ($G$) moduli, from which the Young’s modulus ($E$) and Poisson’s ratio ($\nu$) are further obtained under the isotropic approximation (see Sec. II A of the SM for more details [132]). In solids, sound waves can be generated by either compressional (volumetric) or shear deformation, and such two deformation modes correspond to the pressure (longitudinal) and shear (transverse) waves, respectively. As a consequence, the speed of sound in the same solid material can be different, depending on the specific deformation mode. With the knowledge of elastic moduli, we also calculate the longitudinal ($v_{\rm l}$) and transverse velocities ($v_{\rm t}$), where the detailed formula can be found in Sec. III A of the SM [132]. Overall, our results for the mechanical and elastic properties of silicon are in good agreement with both the finite-difference calculations and experimental measurements.
@@ IV.3 High-symmetry materials
Table 3: The calculated elastic constants, bulk ($K$), shear ($G$) and Young’s ($E$) moduli [GPa], Poisson’s ratio ($\nu$) and longitudinal ($v_{\rm l}$) and transverse ($v_{\rm t}$) sound velocities [m/s] of BaTiO3 from the IFCs of a $14\times 14\times 14$ supercell (this work) and the finite difference of strain-stress relation in thermo_pw. Both the clamped-ion and relaxed-ion results are shown. The elastic moduli are estimated using the Voigt-Reuss-Hill scheme of polycrystalline average.
@@
IV.4 Low-symmetry materials
Now we apply our approach to study the elasticity of the ferroelectric rhombohedral barium titanate (BaTiO3) and the related lattice-dynamical properties. Upon cooling, BaTiO3 displays three consecutive ferroelectric phase transitions: it has the perovskite cubic structure in the high-temperature paraelectric phase, and first undergoes a cubic-to-tetragonal transition around 393 K; the tetragonal ferroelectric phase is then stable until 278 K, where it further transforms into the orthorhombic phase, and below 183 K, BaTiO3 is finally stablized in the rhombohedral phase [137, 138]. In this work, we focus on the low-temperature rhombohedral phase of BaTiO3, with the space group of R3m. Our optimized lattice constant and angle of rhombohedral cell at the LDA level are 3.948 Å and 89.935, respectively, which are close to the measured lattice parameters, 4.004 Å and 89.839 [138]. The phonon dispersions of BaTiO3 are shown in Fig. 3 as a function of the supercell size with increasing level of treatment for the long-range multipolar interactions. As can be seen in Fig. 3(a), the lowest acoustic branch exhibits a imaginary frequency (about -0.1 THz) in the vicinity of the $\Gamma$ point if the standard interpolation is used with only the DD interaction removed. This issue has been attributed to the strong effect of high-order multipole interactions in BaTiO3 [88], such as quadrupole and octupole related terms in the expansion of macroscopic electric field effects. Hence, we have performed a Taylor expansion fit to the unscreened charge density response and the dielectric screening function due to a phonon perturbation up to $\mathcal{O}(q^{2})$ accuracy, and the obtained dynamical multipole tensors and dielectric tensors are summarized in Table 6. Due to the relatively low symmetry of the rhombohedral BaTiO3, the number of independent components for quadrupole are 4, 4 and 10, respectively for Ba, Ti and O atoms, while their corresponding octupole tensors have 6, 6 and 16 symmetry-distinct components. It should be reiterated that the dynamical matrix of BaTiO3 has been found to be non-hermitian because of the unsymmetric multipole tensor regarding to the permutation involving the direction of phonon perturbation, which is a known issue of low-symmetry infrared-active solids [99, 100, 101, 102, 103]. For instance, the second-rank Born effective charge tensor of the O atom is not a symmetric matrix, i.e. $Z_{y}^{z}\neq Z_{z}^{y}$.
[…]
@@ IV.4 Low-symmetry materials
Last but not least, we compute the 6 independent components of the clamped-ion elastic constants of rhombohedral BaTiO3 and present their convergence rate and values in Fig. 4 with increasing level of treatment for the long-range multipolar interactions. We can see that perfect agreement and fast convergence ($10\times 10\times 10$) when subtracting all the long-range between our result and the reference finite-difference results using thermo_pw. We also find that the QQ term has little impact in that material which is in sharp contrast to the results in GaAs where the long-range QQ term plays a vital role in the determination of the correct short-circuit elastic tensor. The lattice-mediated part, Eq. (31), can also be computed with both methods and is reported in Table 3 where some difference can be observed. Indeed, when estimating the contribution of lattice relaxation to the elastic tensor, the structure optimization under a small strain perturbation in DFT simulations does not always reach the global minimum of the Born-Oppenheimer potential energy surface, which gives rise to a difference in the calculated elastic tensor within the finite-difference implementation. Such an issue is a limitation of studying elastic properties with finite difference of strain-stress or strain-energy relation [140] while our approach is able to include the lattice-mediated contributions exactly by using the analytic expression in Eq. (31). Nonetheless, the converged results from the $14\times 14\times 14$ supercell are in relative agreement with the finite-difference elastic constants, where the largest absolute difference is 0.69 GPa and the largest relative difference is 1.24%, as listed in Table 3. The effective elastic moduli and sound velocities of BaTiO3 are also given in Table. 3, and to our best knowledge the experimental measurements of its elastic properties are still unavailable.
@@ IV.5 Elastic tensor and bending rigidity in semi-metallic graphene
Building on the success of our approach for predicting the elastic tensor of bulk solids, we extend it for 2D solids and also compute their bending rigidities. We start with graphene as the prototype 2D material, whose lattice constant is 2.466 Å from our DFT simulation. The calculated phonon dispersion and independent components ($C_{11}$ and $C_{12}$) of the elastic tensor as a function of the supercell size are shown in Fig. 5. Since graphene is a semimetal with a Fermi surface formed by six Dirac points, the long-range effect has been suppressed by the metallic-like screening, and the separate treatment for multipolar interactions is thus not needed. The phonon dispersions and elastic constants of graphene both converge from a 6$\times$6$\times$1 supercell. Note however that two branches are still slightly changing for a larger supercell size due to the well-known Kohn anomalies observed at $\Gamma-E_{2g}$ and $\mathrm{K}-A^{\prime}_{1}$ modes [141, 142]. For our largest 16$\times$16$\times$1 supercell, we find that $C_{11}$ = 352.42 N/m and $C_{12}$ = 64.15 N/m, in good agreement with finite-difference calculations from Ref. [143] and with experiment, see Table 4. We also note that lattice relaxation induced by the internal strain field only has a small effect and the clamped-ion results for $C_{11}$ and $C_{12}$ are found to be 360.19 and 56.38 N/m, respectively.
@@
We then proceed to investigate the bending rigidities of graphene, which are also calculated based on the real-space IFCs but using the higher-order perturbation expansion presented in Sec. III.2. As shown in Fig. 5(b), using the same Voigt notation, the fourth-rank bending rigidity tensor of graphene only has two independent components, $D_{11}$ and $D_{12}$. Unlike the elastic tensor entering the second-order long-wavelength equation of acoustic vibrations, the bending rigidity tensor describes the fourth-order equations of motion for flexural modes, which is more sensitive to the accuracy of IFCs. Furthermore, the bending rigidities of 2D materials can be inferred from the quadratic coefficients of the long-wavelength dispersion of ZA modes, which have a parabolic behavior near the zone center. As can be seen from the inset of Fig. 5(a), the parabolic shape of ZA phonon remains unchanged when varying the supercell size from 4$\times$4$\times$1 to 16$\times$16$\times$1, which is consistent with the observation of a nearly constant bending rigidities of graphene throughout the entire region of supercell size. The obtained $D_{11}$ and $D_{12}$ based on a 16$\times$16$\times$1 supercell are 1.53 and 0.51 eV, respectively.
The principal bending rigidity of graphene $D^{\mathrm{P}}=D_{11}=D_{22}$ was computed previously using Helfrich Hamiltonian model for membranes [79, 80] which gave a value of 1.44 eV [76], in good agreement with our result. In addition, using the Helfrich Hamiltonian method within the cyclic DFT formalism [77, 78], the authors of Ref. [83] obtained a principal bending rigidity of 1.50 eV. Besides, experimental measurement yielded a bending stiffness of monolayer graphene between 1.2 and 1.7 eV [51], further corroborating the accuracy and reliability of our long-wavelength perturbation method. However, we report for the first time the full fourth-rank bending rigidity, which further gives $D_{12}$ = 0.51 eV. This additional information can enable the exploration of bending anisotropy in 2D materials as well as building authentic microscopic to macroscopic models to understand the bending of nanoplates and biological membranes.
@@
Figure 5: (a) Phonon dispersion and (b) independent components of the elastic and bending rigidity tensors in graphene as a function of supercell size. The inset of panel (a) shows the zoom-in acoustic branches. The dotted lines in the panel (b) represent the theoretical elastic constants $C_{11}$/$C_{12}$ and bending rigidity $D_{11}$ taken from Ref. [143] and Ref. [83], respectively.
@@
Table 4: Theoretical and experimental elastic constants, layer ($K$), shear ($G$) and Young’s ($E$) moduli [N/m], Poisson’s ratio ($\nu$) and bending rigidities [eV] of 2D graphene (C), h-BN, MoS2 and InSe. We used a $16\times 16\times 1$ supercell to obtain the elastic and bending rigidity tensors of graphene and h-BN, a $12\times 12\times 1$ supercell is used for MoS2, and a $10\times 10\times 1$ supercell is used for InSe. To obtain the elastic moduli, the 2D polycrystalline average based on the Voigt-Reuss-Hill scheme is utilized. For isotropic 2D materials, the principal bending rigidity is $D^{\mathrm{P}}=D_{11}=D_{22}$, and the Gaussian bending rigidity is calculated as $D^{\mathrm{G}}=-2D_{66}$.
@@
theory 49.38 [155] 13.72 [155] 17.5 [83] 31.55 [155] 17.83 [155] 45.57 [155] 0.28 [155]
The tensorial nature of our calculated bending rigidities allows for a direct evaluation of the Gaussian bending rigidity or Gaussian modulus $D^{\mathrm{G}}$. This quantity describes non-pure bending such as twisting which is dominated by the Gaussian curvature defined as the product of two principal curvatures. In the case of isotropic graphene, $D^{\mathrm{G}}=-2D_{66}$ as derived in Sec. III B of the SM [132] is used to calculate the Gaussian modulus. We obtain a Gaussian modulus of -1.02 eV, in reasonable agreement with the value of -0.70 eV [73] from the Helfrich Hamiltonian method within the density functional tight-binding calculation, while another DFT calculation using the same Helfrich Hamiltonian method gave a value of -1.52 eV [76]. We also notice that a direct measurement of Gaussian modulus is challenging and was only available a few years ago, where atomic force microscopy was used to extract the bending rigidities from a quadratic relationship of adhesion energy between the rolled and unrolled monolayer graphene [145], which gave -1.82 and -1.93 eV for a roll-up nanotube and a bent nanoribbon, respectively. This experimental result is rather different from our computed value. Interestingly, we find that there is no lattice relaxation effect on the bending rigidities of graphene, which results in an additional symmetry relation $D_{12}=D_{66}$. Since $D_{11}-D_{12}=2D_{66}$ holds in hexagonal lattices [156], this means $D_{11}=3D_{66}$ and hence $D^{\mathrm{G}}=-2D^{\mathrm{P}}/3$. In our computational case, such relationship is verified exactly but this is not the case for the current experimental values, therefore awaiting further experimental verification in the future.

### v2

@@ IV.1 Dipole-free materials
Our first example is silicon, which is a simple covalent and infrared-inactive solid. The optimized lattice constant of its diamond-like structure from DFT is 5.469 Å, in close agreement with the experimental value 5.431 Å [141]. Fig. 1 (a) shows the phonon dispersion of silicon with increasing supercell size of IFCs from $4\times 4\times 4$ to $14\times 14\times 14$. Although there is no LO-TO splitting due to its infrared-inactive nature and vanishing Born effective charges, the phonon dispersion only converges from a supercell size of $8\times 8\times 8$. As demonstrated in the inset of Fig. 1 (a), this situation is more severe for long-wavelength acoustic branches around the Brillouin zone center, which is closely linked to the macroscopic elastic properties. The lowest order in the multipole expansion is the quadrupole term, and the QQ term is the only long-range contribution to the dynamical matrix up to the second order in $\mathbf{q}$ in the expansion of Eq. (42). The dynamical quadrupole tensor of silicon can be expressed as $Q_{\kappa\alpha}^{\beta\gamma}=(-1)^{\kappa+1}Q|\varepsilon_{\alpha\beta\gamma}|$ with $\varepsilon_{\alpha\beta\gamma}$ being the Levi-Civita symbol, so it only has one independent component when $\alpha\neq\beta\neq\gamma$. Our fitting procedure based on the Taylor expansion of unscreened charge density response in Eq. (43) gives a value of 13.76 $e~\rm{bohr}$ for $|Q_{x}^{yz}|$ which agrees well with the previous calculations either from DFPT (13.67 $e~\rm{bohr}$ [136]) or a fit to the analytic long-range electron-phonon matrix elements (13.66 $e~\rm{bohr}$ [142]). However, our further interpolation of dynamical matrix with the QQ term taken into account shows an almost unchanged phonon dispersion for silicon, except for the case using a $4\times 4\times 4$ supercell which is too small to ensure the good decay of IFCs in real space. Therefore, the inclusion of quadrupole term does not help to improve the convergence rate of lattice-dynamical properties for silicon, and only the phonon dispersion based on the standard interpolation scheme is illustrated in Fig. 1 (a). We report the dynamical multipole tensors and dielectric properties of silicon in Table 1, although the octupole tensor is not used here due to the vanishing DO interactions in silicon.
@@
Silicon has the highest cubic symmetry and there is only three independent components ($C_{11}$, $C_{12}$ and $C_{44}$) for its elastic tensor (see Sec. II of the SM [118] for a brief introduction of Voigt notation). The calculated elastic stiffness tensor based on the real space IFCs and long-wavelength expansion as a function of supercell size is shown in Fig. 1 (b). Similar to the results of phonon dispersion, the elastic constants of silicon also begin to converge after a supercell of $8\times 8\times 8$ with the relative error of $4.3\%$ and $2.9\%$ to the reference values from finite-difference calculations for $C_{12}$ and $C_{44}$, respectively, while the $C_{11}$ component has already converged after a $6\times 6\times 6$ phonon grid. The results displayed here are calculated based on the total IFCs as silicon is infrared-inactive, without the removal of any high-order multipolar interaction. A further subtraction of the QQ term in the dynamical matrix of silicon does not modify its elastic tensor, which implies that the quadrupole will not couple with the mechanical field therein; this observation is in agreement with the fact that silicon is not a piezoelectric crystal because of the cubic symmetry and the existence of spatial inversion. Using the obtained quadrupole values, the calculated piezoelectric tensor via Martin’s formula [114] is indeed zero. In addition, the detailed comparison of the elastic constants, moduli and sound velocities of silicon calculated from our method based on a supercell of $14\times 14\times 14$ and the finite difference implemented in the thermo_pw code [59] as well as the experimental values [137, 138] are listed in Table 2. The Voigt-Reuss-Hill (VRH) scheme of polycrystalline average is adopted to calculate the bulk ($K$) and shear ($G$) moduli, from which the Young’s modulus ($E$) and Poisson’s ratio ($\nu$) are further obtained under the isotropic approximation (see Sec. II A of the SM [118] for more details). In solids, sound waves can be generated by either compressional (volumetric) or shear deformation, and two such deformation modes correspond to the pressure (longitudinal) and shear (transverse) waves, respectively. As a consequence, the speed of sound in the same solid material can be different, depending on the specific deformation mode. With the knowledge of elastic moduli, we also calculate the longitudinal ($v_{\rm l}$) and transverse velocities ($v_{\rm t}$), where the detailed formula can be found in Sec. III A of the SM [118]. Overall, our results for the mechanical and elastic properties of silicon are in good agreement with both the finite-difference calculations and experimental measurements.
@@ IV.3 High-symmetry materials
[…]
IV.4 Low-symmetry materials
Now we apply our approach to study the elasticity of the ferroelectric rhombohedral barium titanate (BaTiO3) and the related lattice-dynamical properties. Upon cooling, BaTiO3 displays three consecutive ferroelectric phase transitions: it has the perovskite cubic structure in the high-temperature paraelectric phase, and first undergoes a cubic-to-tetragonal transition around 393 K; the tetragonal ferroelectric phase is then stable until 278 K, where it further transforms into the orthorhombic phase, and below 183 K, BaTiO3 is finally stablized in the rhombohedral phase [146, 147]. In this work, we focus on the low-temperature rhombohedral phase of BaTiO3, with the space group of R3m. Our optimized lattice constant and angle of rhombohedral cell at the LDA level are 3.948 Å and 89.935, respectively, which are close to the measured lattice parameters, 4.004 Å and 89.839 [147]. The phonon dispersions of BaTiO3 are shown in Fig. 3 as a function of the supercell size with increasing level of treatment for the long-range multipolar interactions. As can be seen in Fig. 3(a), the lowest acoustic branch exhibits an imaginary frequency (about -0.1 THz) in the vicinity of the $\Gamma$ point if the standard interpolation is used with only the DD interaction removed. This issue has been attributed to the strong effect of high-order multipole interactions in BaTiO3 [95], such as quadrupole- and octupole-related terms in the expansion of macroscopic electric field effects. Hence, we have performed a Taylor expansion fit to the unscreened charge density response and the dielectric screening function due to a phonon perturbation up to $\mathcal{O}(q^{2})$ accuracy, and the obtained dynamical multipole tensors and dielectric tensors are summarized in Table 6. Due to the relatively low symmetry of the rhombohedral BaTiO3, the number of independent components for quadrupole are 4, 4, and 10, respectively, for Ba, Ti and O atoms, while their corresponding octupole tensors have 6, 6, and 16 symmetry-distinct components. It should be reiterated that the dynamical matrix of BaTiO3 has been found to be non-hermitian because of the unsymmetric multipole tensor regarding to the permutation involving the direction of phonon perturbation, which is a known issue of low-symmetry infrared-active solids [106, 107, 108, 109, 110]. For instance, the second-rank Born effective charge tensor of the O atom is not a symmetric matrix, i.e. $Z_{y}^{z}\neq Z_{z}^{y}$.
Table 3: The calculated elastic constants ($C_{ij}$), bulk ($K$), shear ($G$) and Young’s ($E$) moduli [GPa], Poisson’s ratio ($\nu$) and longitudinal ($v_{\rm l}$) and transverse ($v_{\rm t}$) sound velocities [m/s] of BaTiO3 from the IFCs of a $14\times 14\times 14$ supercell (this work) and the finite difference of strain-stress relation in thermo_pw. Both the clamped-ion and relaxed-ion results are shown. The elastic moduli are estimated using the Voigt-Reuss-Hill scheme of polycrystalline average.
@@ IV.4 Low-symmetry materials
Last but not least, we compute the six independent components of the clamped-ion elastic constants of rhombohedral BaTiO3 and present their convergence rate and values in Fig. 4 with increasing level of treatment for the long-range multipolar interactions. We can see that perfect agreement and fast convergence ($10\times 10\times 10$) when subtracting all the long-range between our result and the reference finite-difference results using thermo_pw. We also find that the QQ term has little impact in that material, which is in sharp contrast to the results in GaAs where the long-range QQ term plays a vital role in the determination of the correct short-circuit elastic tensor. The lattice-mediated part, Eq. (31), can also be computed with both methods and is reported in Table 3 where some difference can be observed. Indeed, when estimating the contribution of lattice relaxation to the elastic tensor, the structure optimization under a small strain perturbation in DFT simulations does not always reach the global minimum of the Born-Oppenheimer potential energy surface, which gives rise to a difference in the calculated elastic tensor within the finite-difference implementation. Such an issue is a limitation of studying elastic properties with finite difference of strain-stress or strain-energy relation [71], while our approach is able to include the lattice-mediated contributions exactly by using the analytic expression in Eq. (31). Nonetheless, the converged results from the $14\times 14\times 14$ supercell are in relative agreement with the finite-difference elastic constants, where the largest absolute difference is 0.69 GPa and the largest relative difference is 1.24%, as listed in Table 3. The effective elastic moduli and sound velocities of BaTiO3 are also given in Table. 3, and to our best knowledge the experimental measurements of its elastic properties are still unavailable.
@@ IV.5 Elastic tensor and bending rigidity in semi-metallic graphene
Building on the success of our approach for predicting the elastic tensor of bulk solids, we extend it for 2D solids and also compute their bending rigidities. We start with graphene as the prototype 2D material, whose lattice constant is 2.466 Å from our DFT simulation. The calculated phonon dispersion and independent components ($C_{11}$ and $C_{12}$) of the elastic tensor as a function of the supercell size are shown in Fig. 5. Since graphene is a semimetal with a Fermi surface formed by six Dirac points, the long-range effect has been suppressed by the metallic-like screening, and the separate treatment for multipolar interactions is thus not needed. The phonon dispersions and elastic constants of graphene both converge from a $\times6$$\times$1 supercell. However, the uppermost two optical branches are still slightly changing for a larger supercell size due to the well-known Kohn anomalies observed at $\Gamma-E_{2g}$ and $\mathrm{K}-A^{\prime}_{1}$ modes [149, 150]. For our largest $\times16$$\times$1 supercell, we find that $C_{11}$ = 352.42 N/m and $C_{12}$ = 64.15 N/m, in good agreement with finite-difference calculations from Ref. [151] and with experiment, see Table 4. We also note that lattice relaxation induced by the internal strain field only has a small effect and the clamped-ion results for $C_{11}$ and $C_{12}$ are found to be 360.19 and 56.38 N/m, respectively.
@@
We then proceed to investigate the bending rigidities of graphene, which are also calculated based on the real-space IFCs but using the higher-order perturbation expansion presented in Sec. III.2. As shown in Fig. 5(b), using the same Voigt notation, the fourth-rank bending rigidity tensor of graphene only has two independent components, $D_{11}$ and $D_{12}$. Unlike the elastic tensor entering the second-order long-wavelength equation of acoustic vibrations, the bending rigidity tensor describes the fourth-order equations of motion for flexural modes, which is more sensitive to the accuracy of IFCs. Furthermore, the bending rigidities of 2D materials can be inferred from the quadratic coefficients of the long-wavelength dispersion of ZA modes, which have a parabolic behavior near the zone center. As can be seen from the inset of Fig. 5(a), the parabolic shape of ZA phonon remains unchanged when varying the supercell size from $\times4$$\times$1 to $\times16$$\times$1, which is consistent with the observation of a nearly constant bending rigidities of graphene throughout the entire region of supercell size. The obtained $D_{11}$ and $D_{12}$ based on a $\times16$$\times$1 supercell are 1.53 and 0.51 eV, respectively. The principal bending rigidity of graphene $D^{\mathrm{P}}=D_{11}=D_{22}$ was computed previously using the Helfrich Hamiltonian model for membranes [85, 86], which gave a value of 1.44 eV [82], in good agreement with our result. In addition, using the Helfrich Hamiltonian method within the cyclic DFT formalism [83, 84], the authors of Ref. [89] obtained a principal bending rigidity of 1.50 eV. Besides, experimental measurement yielded a bending stiffness of monolayer graphene between 1.2 and 1.7 eV [55], further corroborating the accuracy and reliability of our long-wavelength perturbation method.
@@
Figure 5: (a) Phonon dispersion and (b) independent components of the elastic and bending rigidity tensors in graphene as a function of supercell size. The inset of panel (a) shows the enlarged acoustic branches. The dotted lines in the panel (b) represent the theoretical elastic constants $C_{11}$$C_{12}$ and bending rigidity $D_{11}$ taken from Ref. [151] and Ref. [89], respectively.
The tensorial nature of our bending rigidity formalism also allows for a direct evaluation of the Gaussian bending rigidity or Gaussian modulus $D^{\mathrm{G}}$. This quantity describes non-pure bending such as twisting which is dominated by the Gaussian curvature defined as the product of two principal curvatures. In the case of an isotropic system, $D^{\mathrm{G}}=-2D_{66}$ as derived in Sec. III B of the SM [118] is used to calculate the Gaussian modulus. We obtain a Gaussian modulus of -1.02 eV, in reasonable agreement with the value of -0.70 eV [79] from the Helfrich Hamiltonian method within the density functional tight-binding calculation, while another DFT calculation using the same Helfrich Hamiltonian method gave a value of -1.52 eV [82]. We also notice that a direct measurement of Gaussian modulus is challenging and was only available a few years ago, where atomic force microscopy was used to extract the bending rigidities from a quadratic relationship of adhesion energy between the rolled and unrolled monolayer graphene [152], which gave -1.82 and -1.93 eV for a roll-up nanotube and a bent nanoribbon, respectively. This experimental result is rather different from our computed value. Interestingly, we find that there is no lattice relaxation effect on the bending rigidities of graphene, which results in an additional symmetry relation $D_{12}=D_{66}$. Since $D_{11}-D_{12}=2D_{66}$ holds in hexagonal lattices [153], this means $D_{11}=3D_{66}$ and hence $D^{\mathrm{G}}=-2D^{\mathrm{P}}/3$ in the case of graphene. In our computational case, such relationship is verified exactly but this is not the situation for the current experimental values, therefore awaiting further experimental verification in the future. One possibility for the observed discrepancy in the Gaussian modulus of graphene compared to those from the Helfrich Hamiltonian method is the existence of local geometric distortions due to the use of a relatively large curvature in their calculations, i.e. not in the linear response region. In addition, the calculations using fullerene-like structures require an additional energy correction due to the existence of 12 pentagon rings, which causes a poor fitting performance for the Gaussian modulus [82]. Instead, our perturbative approach uses only a single unit cell and is exact in the infinitely small curvature limit. Moreover, the probe-induced topological defects and their attractive interactions could explain the more negative Gaussian modulus obtained in experiments [152]. Importantly, the obtained complete bending rigidity tensor of a general monolayer can enable the exploration of bending anisotropy as well as building authentic microscopic to macroscopic models to understand the bending of nanoplates and biological membranes.
@@
Table 4: Theoretical and experimental elastic constants ($C_{ij}$), layer ($K$), shear ($G$) and Young’s ($E$) moduli [N/m], Poisson’s ratio ($\nu$) and bending rigidities ($D_{ij}$) [eV] of 2D graphene (C), h-BN, MoS2, and InSe. We used a $16\times 16\times 1$ supercell to obtain the elastic and bending rigidity tensors of graphene and h-BN, a $12\times 12\times 1$ supercell is used for MoS2, and a $10\times 10\times 1$ supercell is used for InSe. To obtain the elastic moduli, the 2D polycrystalline average based on the Voigt-Reuss-Hill scheme is utilized. For isotropic 2D materials, the principal bending rigidity is $D^{\mathrm{P}}=D_{11}=D_{22}$, and the Gaussian bending rigidity is calculated as $D^{\mathrm{G}}=-2D_{66}$. We note that the layer modulus is the 2D equivalent of the bulk modulus and the experiment data are all assumed to be obtained at room temperature including those specifically reported as such.
@@
theory 49.38 [164] 13.72 [164] 15.3, 17.5 [89] 31.55 [164] 17.83 [164] 45.57 [164] 0.28 [164]